# Proper Body Landmark Subset Enables More Accurate and 5X Faster Recognition of Isolated Signs in LIBRAS

### v2

@@ 1 Introduction
Figure 1: Pipeline of the proposed landmark-based approach for isolated sign recognition in LIBRAS. Video frames are processed with MediaPipe to extract landmarks, which are selected (subset selection), interpolated (imputation), encoded as 2-D skeleton images, and classified by a CNN to predict the sign label.
Artificial intelligence has enabled significant advancements in accessible technologies across various domains (Hossain et al., , 2023; Pereira et al., , 2025; Pantera et al., , 2025). In particular, there is a growing demand for solutions that support individuals with partial or total hearing loss, a population estimated at $1.57$ billion in 2019, and projected to reach $2.45$ billion worldwide by 2050 (Guo et al., , 2024; Nieman et al., , 2024).
@@
In this context, several works have addressed sentence-level (continuous) (Wang et al., , 2025; Hu et al., , 2023) and word-level (isolated) (Alves et al., , 2024; Kumari and Anand, , 2024) recognition of Sign Language (SL) from video sequences. In the latter, which is the focus of our investigation, the objective is to identify the sign (“word”) performed by an individual (signer) recorded in front of a camera, assuming that a single sign is performed. This application is particularly helpful for self-paced SL training platforms (Starner et al., , 2024), visual keyword search for SL (Tamer and Saraçlar, , 2020), dictionary lookup by example (Bohacek and Hassan, , 2023), as well as query-by-example sign spotting to locate occurrences in continuous videos (Varol et al., , 2022; Vázquez Enríquez et al., , 2022), and cross-modal retrieval to search sign videos from free-form text (Duarte et al., , 2022).
Most works addressing Isolated Sign Language Recognition (ISLR) from video sequences leverage deep neural networks, such as 2- and 3-D Convolutional Neural Networks (CNNs) (Alves et al., , 2024; De Castro et al., , 2023), Gated Recurrent Units (GRUs) (Shen et al., , 2024), Long Short-term Memory (LSTM) networks (Aly and Aly, , 2020; Rastgoo et al., , 2021), and Transformers (Ertürk et al., , 2025; Boháček and Hrúz, , 2022; Pu et al., , 2024). Promising results have been achieved by combining multiple modalities. For instance, Rastgoo et al., (2021) employed a 2-D CNN to extract four types of spatial features: global frame features, local hand features, optical flow, and hand-pose features. These features were fused with depth maps generated by Microsoft Kinect, and subsequently fed into an LSTM for temporal processing.
@@
Multimodality was also explored by De Castro et al., (2023), who employed 3-D CNNs across different scenarios, ranging from the use of raw RGB data to the integration of optical flow, synthetically generated depth maps, segmented hands and faces, and distance/velocity maps derived from body landmarks (joints and other keypoints). Their best results were achieved with a complex multi-stream architecture in which each branch processed: (i) joint distance and velocity maps; (ii) segmented hands; (iii) segmented faces; and (iv) the original RGB frames augmented with a depth map. Despite its effectiveness, the approach introduces significant complexity and computational overhead, as landmark extraction depends on OpenPose (Cao et al., , 2021), a framework that delivers good accuracy but at a high computational cost.
To address the complexity of the multi-stream design, Alves et al., (2024) proposed a single-stream pipeline where body landmarks – covering the trunk, hands, and face – are extracted from RGB frames and transformed into a 2-D skeleton-image representation (Memmesheimer et al., , 2022), which is then classified by a 2-D CNN. Their approach outperformed De Castro et al., (2023), achieving state-of-the-art recognition accuracy on MINDS-Libras (Rezende et al., , 2021) and LIBRAS-UFOP (Cerna et al., , 2021), the most widely used LIBRAS benchmarks in the context of ISLR. However, their method still relies on OpenPose for landmark extraction, incurring considerable computational cost and posing a barrier to real-time applications.
This limitation has motivated the exploration of lightweight frameworks for landmark extraction, such as MediaPipe (Lugaresi et al., , 2019), enabling greater efficiency and scalability. Ertürk et al., (2025) used MediaPipe to extract landmarks, which are processed by a Transformer network (Vaswani et al., , 2017), achieving over 90% accuracy in Turkish SL. Similarly, Laines et al., (2023) convert MediaPipe holistic landmarks (hands, face, and upper body) into Tree-Structure Skeleton Images and report state-of-the-art results among skeleton-based models on WLASL, with competitive performance on AUTSL. Luna-Jiménez et al., (2023) employed a Transformer network over MediaPipe landmarks (hands + pose). They showed that using MediaPipe improves WLASL100 accuracy by 13.18 percentage points compared to non-MediaPipe features (Luna-Jiménez et al., , 2023).
@@
The time efficiency enabled by MediaPipe has motivated us to investigate its impact on the ISLR framework proposed by Alves et al., (2024). However, preliminary exploration showed that using all landmark points provided by MediaPipe dramatically reduced recognition accuracy. Based on this observation, we conducted this investigation to verify whether landmark subsets used in the literature, as well as in winning solutions publicly available from SLR challenges, enable the use of MediaPipe for MINDS-Libras and LIBRAS-UFOP. The main findings in this paper are:
@@
  • Selecting an appropriate subset of landmarks substantially improves recognition performance, achieving higher accuracy while reducing input dimensionality.
@@
  • Applying a simple spline-based imputation for missing landmarks – common in MediaPipe detections – further enhances accuracy.
  • The proposed method, using the best-performing landmark strategy, outperforms existing approaches in the literature.
[…]
The remainder of this paper is organized as follows. Section 2 presents the landmark-based framework for recognition of isolated signs. Section 3 describes the experimental methodology, while the results are presented and analyzed in Section 4. Finally, Section 5 discusses the concluding remarks and directions for future work.
@@ 2 Landmark-based ISLR Framework
The landmark-based ISLR pipeline is illustrated in Figure 1. The input sample is a video sequence associated with a single sign. Each RGB frame of the video is processed individually to detect body landmarks. This procedure relies on the MediaPipe Holistic model (Lugaresi et al., , 2019), which combines pose, face, and hand markers into a unified full-body estimation. The detected landmark points are organized into a CSV file, where the columns represent the xx/$y$ normalized coordinates of each specific point, and the rows represent the temporal sequence.
A subset of the detected landmarks (i.e., columns of the CSV file) is selected according to a predefined strategy. Due to noise in the landmark detection process, several missing points can be observed in the CSV file. To address this issue, we applied spline-based interpolation to impute the missing points. This approach shortens detection gaps, yielding temporally continuous trajectories for each point. The processed landmarks are subsequently converted into a 2-D skeleton image representation (Memmesheimer et al., , 2022), which compactly encodes the spatial structure and temporal dynamics of the landmarks. Finally, the image is resized to $224\times 224$ and fed into a CNN (ResNet-18), which outputs the sign label based on the highest probability.
The following subsections detail each stage of the pipeline: landmark extraction, landmark subset selection, imputation, image encoding, and the training process used to deploy the classification model.
@@ 2.1 Landmarks Extraction
Figure 2: Landmarks subset selection: the columns show the set of landmarks utilized in each each strategy and the respective count across the body parts.
The MediaPipe framework has been extensively adopted in real-time human perception tasks (Uddin et al., , 2025; Sultana et al., , 2024; Bagga and Yang, , 2024), due to its ability to capture manual articulators (hands), upper-body kinematics, and orofacial cues. Our ISLR pipeline relies on the Holistic model, which outputs full-body landmarks, including (by default) 543 points (Figure 2 (baseline)): 468 for the face, 33 for the pose (estimating body posture per frame across the entire body), and 21 per hand (42 in total). Note that the pose landmarks include a few coarse points around the head (e.g., nose and eyes), but the detailed facial landmarks are provided as a separate component. MediaPipe also provides depth information ($z$-coordinate) and visibility scores. Our approach, however, operates strictly in 2-D and therefore discards the additional information provided by the MediaPipe framework.
Settings
The minimum detection confidence and minimum tracking thresholds in MediaPipe were set to 0.4. This value proved effective in preliminary tests for stabilizing landmark extraction. The former triggers the initial detection of the articulators (e.g., hands, face, and pose) in each frame, while the latter controls temporal tracking to maintain consistent trajectories across frames.
@@ 2.2 Landmarks Subset Selection
As mentioned, 468 out of the 543 landmarks provided by the Holistic model correspond to the face alone, potentially introducing redundancy (non-linguistic variation) and noise. Based on the premise that appropriate landmark subsets are key to improving recognition accuracy, five strategies for landmark selection are considered in this study: the full set of points (baseline) and four subsets focusing on the most relevant articulators for SLR. The landmark (sub)sets are described in the following.
Baseline (All)
This strategy (Figure 2, col. 1) uses all 543 landmarks provided by MediaPipe Holistic, serving as a reference for comparison against the other strategies evaluated in this work. By retaining the full configuration, it maximizes spatial and articulatory information, preserving facial micro-configurations (mouth, eyebrows, and gaze), upper-limb kinematics, and manual gestures. The trade-off is a substantially larger input data and increased susceptibility to noise.
Laines et al., (2023)
This strategy (Figure 2, col. 2) employs 68 landmarks distributed across the face, pose, and hands, prioritizing regions functionally relevant for signing (lips/expressivity, shoulder girdle and upper-limb joints, plus key points on both hands). As noted by the authors, the goal is to reduce dimensionality while preserving the articulatory mechanisms most critical for ISLR: manual configurations, arm/forearm orientation and perspective, and facial cues related to the mouth and expressiveness. Compared to the baseline (All), this approach is expected to provide greater robustness to noise and lower computational cost, with potential performance gains when signs are strongly driven by the hands and salient facial/head regions.
Arcanjo et al., (2024)
@@
ASL Signs Challenge 1st-place (ASL-1st)
This strategy (Figure 2, col. 4) adopts the subset from the winning solution (Sohn, , 2023) of the Google ASL Signs Challenge on Kaggle, which focuses on isolated signs. A total of 118 landmarks were selected from the face and hands, excluding body pose. The emphasis is on manual configurations and the orofacial region (lips/jaw), which are distinctive in many lexicons, while global pose is discarded to reduce dimensionality and dependence on posture. The underlying observation is that, for many isolated signs, the differential cues lie in hand shape, contact/relative position to the face, and lip articulation rather than in trunk alignment.
ASL Signs Challenge 2nd-place (ASL-2nd)
Also derived from the Google ASL Signs Challenge on Kaggle, this configuration (Figure 2, col. 5) replicates the subset from the runner-up solution (Toporov et al., , 2023), comprising 80 keypoints concentrated on the lips, hands, and body pose. Unlike ASL-1st, it preserves a core of pose landmarks to retain postural context (shoulder, elbow, and wrist angles) while still prioritizing manual articulation and the labial region. The aim is to balance fine-grained hand discrimination with sufficient global information to disambiguate signs with distinct trajectories or ranges.
@@ 2.3 Spline-Based Landmark Imputation
The imputation procedure addresses detection gaps resulting from MediaPipe failures. To this end, we apply piecewise spline interpolation, treating each landmark-coordinate combination across time as an independent function to be reconstructed. Missing values are estimated from their short temporal neighborhoods (window size of 5), leveraging information from both past and future frames. Specifically, we employ a cubic spline for robust reconstruction when at least four points are available in the series; otherwise, the procedure defaults to linear interpolation. The short-window constraint preserves kinematic plausibility and prevents extrapolation beyond the observed range. This step stabilizes the spatiotemporal trajectories prior to the encoding stage, mitigating discontinuities that could otherwise degrade classifier training.
2.4 Image Encoding
Landmarks are encoded as images using SkeletonDML, an algorithm proposed by Memmesheimer et al., (2022), originally applied to represent skeletal joints in few-shot action recognition. To the best of our knowledge, Alves et al., (2024) were the first to apply Skeleton-DML to ISLR, obtaining promising results on LIBRAS datasets. The underlying idea of such representation applies broadly to space–time coordinates, such as general body landmarks. Adapting it to the 2-D case is straightforward by disregarding depth ($z$-coordinate), which simply reduces the encoded data volume.
Given $L$ landmarks and $T$ frames, the 2-D version of the Skeleton-DML encoding procedure consists of creating two $L\times T$ matrices (tensors), one for each of the $x$ and $y$ coordinates. These tensors are reshaped so that three consecutive column entries are organized as channels, forming a 3-channel image. The reshaped tensors are then horizontally concatenated, forming a single image. More details are provided in the original work (Memmesheimer et al., , 2022). For a more formal description, the reader is referred to Alves et al., (2024).
@@ 2.5 Training Protocol
The training protocol follows the procedure utilized by Alves et al., (2024), with a single modification, that we train for 30 epochs instead of the original 20. At every epoch, we perform on-the-fly data augmentation directly on the landmark sequences prior to their encoding into 2-D images, consisting of rotation, zoom, translation, and horizontal flip to increase data diversity without changing the number of samples.
The classifier model is a ResNet-18 initialized with ImageNet weights. Its final convolutional output is first passed through a global average pooling layer, producing a 512-dimensional feature vector. This vector is then fed into two fully connected layers: a 128-unit layer with ReLU activation, followed by an output layer matching the number of classes. The network input is $224\times 224$, while landmark-encoded images have size $126\times(2T/3)$, where $T$ is the number of frames of that particular input video. For the evaluated datasets, $2T/3<224$, so resizing to $224\times 224$ results in upsampling without information loss.
Optimization leverages Adam to minimize cross-entropy, with batch size $64$, learning rate $10^{-4}$, and 30 epochs. Early stopping is triggered if the validation loss fails to decrease for five consecutive epochs. To prevent overfitting, we apply (i) batch normalization after the 512-dimensional layer, (ii) dropout (with 50% probability) after the 128-unit layer, and (iii) L2 regularization (weight decay of $10^{-4}$) on all trainable parameters. The model selected for evaluation is the checkpoint achieving the highest validation accuracy.
@@ 3 Experimental Methodology
This section outlines the experimental framework used for performance assessment, including the Brazilian Sign Language (LIBRAS) datasets, evaluation metrics, experiments, and the computational setup.
@@ 3.1 Datasets
Two public popular ISLR LIBRAS datasets were considered in this investigation: MINDS-Libras (Rezende et al., , 2021) and LIBRAS-UFOP (Cerna et al., , 2021). In particular, these collections enable direct comparison with Alves et al., (2024), whose methodology motivated this study.

### v3

@@ Abstract
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Artificial intelligence has enabled significant advancements in accessible technologies across various domains, in particular, to support individuals with partial or total hearing loss: a population estimated at $1.57$ billion in 2019, and projected to reach $2.45$ billion worldwide by 2050 [9]. Literature has addressed sentence-level (continuous) [10] and word-level (isolated) [1] recognition of Sign Language (SL) from video sequences. In the latter, the focus of this work, the goal is to identify a single sign (“word”) performed by an individual (signer) recorded in front of a camera. This application is particularly helpful for self-paced SL training platforms [21], visual keyword search for SL [22], dictionary lookup by example [3], as well as query-by-example sign spotting to locate occurrences in continuous videos [24], and cross-modal retrieval to search sign videos from free-form text [7].
@@
Most works on Isolated Sign Language Recognition (ISLR) leverage deep neural networks, such as 2- and 3-D Convolutional Neural Networks (CNNs) [1, 6], Gated Recurrent Units (GRUs) [19], Long Short-term Memory (LSTM) networks [17], and Transformers [8]. Promising results have been achieved by combining multiple modalities. For instance, De Castro [6] proposed a complex multi-stream 3-D CNN architecture in which each branch processed: joint distance and velocity maps; segmented hands; segmented faces; and the raw RGB frames augmented with a depth map. Despite its effectiveness, the approach introduces significant complexity and computational overhead, as landmark extraction depends on OpenPose [4], a framework that delivers good accuracy but at a high computational cost.
Conversely, Alves et al. [1] proposed a simpler, yet effective, single-stream pipeline. Body landmarks (trunk, hands, and face) are extracted from RGB frames and transformed into a 2-D skeleton-image representation [14], which is fed into a CNN classifier. They achieved state-of-the-art recognition accuracy on the two most popular LIBRAS ISLR datasets (MINDS-Libras [18] and LIBRAS-UFOP [5]), however, at the cost of obtaining landmarks with the OpenPose framework. To enhance efficiency and scalability, lightweight frameworks, such as MediaPipe [12], have been explored. In [8, 13], the landmarks obtained with MediaPipe are directly processed by a Transformer network [25], while in [11], the MediaPipe holistic landmarks (hands, face, and upper body) are converted into Tree-Structure Skeleton Images, later processed by a 2-D CNN.
@@
The time efficiency enabled by MediaPipe has motivated us to investigate its impact on the ISLR framework proposed by Alves et al. [1]. However, preliminary exploration showed that using all landmarks provided by MediaPipe dramatically reduced recognition accuracy. Based on this observation, we investigate whether landmark subsets used in the literature, as well as in winning solutions publicly available from SLR challenges, enable the use of MediaPipe for MINDS-Libras and LIBRAS-UFOP. The main findings in this paper are:
@@
  • A proper subset of landmarks substantially improves recognition performance while reducing input dimensionality.
@@
  • Simple spline-based imputation for missing landmarks further enhances accuracy.
  • The best-performing landmark subset outperformed the literature on the MINDS-Libras and LIBRAS-UFOP datasets.
  • Leveraging MediaPipe yields a significant speed-up compared to the OpenPose counterpart [1]: $6\times$ for landmark extraction and $5\times$ for the overall pipeline.
Figure 1: Overview of the ISLR framework. MediaPipe is used to detect landmarks, which are selected, interpolated, encoded as 2-D skeleton images, and then classified by a CNN.
@@ 2 Landmark-based ISLR Framework
The ISLR framework is illustrated in Figure 1. The input is a video sequence comprising a single sign. Body landmarks are extracted from each RGB frame using the MediaPipe Holistic model [12], which combines pose, face, and hand markers into a unified full-body estimation. The landmarks are organized into a CSV file, where an even/odd column represents the xx/$y$ normalized coordinate of a particular landmark, whereas the rows represent the temporal sequence. A subset of the detected landmarks is selected according to a predefined strategy. To impute missing points (due to noisy detection), we apply spline-based interpolation, modeling temporally continuous trajectories for each point. The processed landmarks are then converted into a 2-D skeleton image representation [14], which encodes the spatial structure and temporal dynamics of the landmarks. Finally, the image is resized to $224\times 224$ and fed into a 2-D CNN, which outputs the highest-probability sign.
The following sections detail the landmark extraction, landmark subset selection, imputation, image encoding, and the training process to deploy the classification model.
@@ 2.1 Landmarks Extraction
Figure 2: Landmarks subset selection: the columns show the set of landmarks utilized in each strategy and the respective count across the body parts.
MediaPipe has been widely used in real-time human perception tasks due to its ability to capture manual articulators (hands), upper-body kinematics, and orofacial cues. Our ISLR pipeline relies on the Holistic model, which outputs full-body landmarks, including (by default) 543 points (Figure 2 (baseline)): 468 for the face, 33 for the pose (estimating body posture per frame across the entire body), and 21 per hand (42 in total). Note that the pose landmarks include a few coarse points around the head (e.g., nose and eyes), but the detailed facial landmarks are provided as a separate component. MediaPipe also provides depth information ($z$-coordinate) and visibility scores. Our approach, however, operates strictly in 2-D and therefore discards the additional information provided by MediaPipe.
Settings. The minimum detection confidence and minimum tracking thresholds in MediaPipe were set to 0.4, a value proven effective in preliminary tests for stabilizing landmark extraction. The former triggers the initial detection of the articulators (e.g., hands, face, and pose) in each frame, while the latter controls temporal tracking to maintain consistent trajectories across frames.
@@ 2.2 Landmarks Subset Selection
As mentioned, 468 out of the 543 landmarks provided by the Holistic model correspond to the face alone, potentially introducing redundancy (non-linguistic variation) and noise. Based on the premise that appropriate landmark subsets are key to improve recognition, five strategies for landmark selection are considered in this study: the full set of points (baseline) and four subsets focusing on the most relevant articulators for SLR.
2.2.1 Baseline (All)
The baseline (Figure 2, col. 1) uses all 543 landmarks provided by the Holistic model, serving as a reference for comparison against the other strategies. By retaining the full configuration, it maximizes spatial and articulatory information, preserving facial micro-configurations (mouth, eyebrows, and gaze), upper-limb kinematics, and manual gestures. The trade-off is a substantially larger input data and increased susceptibility to noise.
2.2.2 Laines [11]
This strategy (Figure 2, col. 2) employs 68 landmarks distributed across the face, pose, and hands, prioritizing regions functionally relevant for signing (lips/expressivity, shoulder girdle and upper-limb joints, plus key points on both hands). The goal is to reduce dimensionality while preserving the articulatory mechanisms most critical for ISLR: manual configurations, arm/forearm orientation and perspective, and facial cues related to the mouth and expressiveness. Compared to the baseline (All), this strategy is expected to be more robust to noise, with potential performance gains when signs are strongly driven by the hands and salient facial/head regions.
2.2.3 Arcanjo [2]
@@
2.2.4 ASL Signs Challenge 1st-place (ASL-1st)
This strategy (Figure 2, col. 4) adopts the subset from the winning solution [20] of the Google ASL Signs Challenge on Kaggle (focus on isolated signs). A total of 118 landmarks were selected from the face and hands, excluding body pose. The emphasis is on manual configurations and the orofacial region (lips/jaw), which are distinctive in many lexicons, while global pose is discarded to reduce dimensionality and dependence on posture. The underlying observation is that, for many isolated signs, the differential cues lie in hand shape, contact/relative position to the face, and lip articulation rather than in trunk alignment.
2.2.5 ASL Signs Challenge 2nd-place (ASL-2nd)
Also derived from the Google ASL Signs Challenge on Kaggle, this configuration (Figure 2, col. 5) replicates the subset from the runner-up solution [23], comprising 80 keypoints concentrated on the lips, hands, and body pose. Unlike ASL-1st, it preserves a core of pose landmarks to retain postural context (shoulder, elbow, and wrist angles) while still prioritizing manual articulation and the labial region. The aim is to balance fine-grained hand discrimination with sufficient global information to disambiguate signs with distinct trajectories or ranges.
@@ 2.3 Spline-Based Landmark Imputation
Imputation is achieved through piecewise spline interpolation, where each landmark-coordinate time series is treated as an independent function to be reconstructed. Missing values are estimated from their short temporal neighborhoods (window size of 5), leveraging information from both past and future frames. Specifically, we employ a cubic spline for robust reconstruction when at least four points are available in the series; otherwise, the procedure defaults to linear interpolation. The short-window constraint preserves kinematic plausibility and prevents extrapolation beyond the observed range.
2.4 Image Encoding via Skeleton-DML
Landmarks are encoded as images using Skeleton-DML [14], originally applied to represent skeletal joints in few-shot action recognition. To the best of our knowledge, Alves et al. [1] were the first to apply Skeleton-DML to ISLR, obtaining promising results on LIBRAS datasets. The underlying idea of such representation applies broadly to space–time coordinates, such as general body landmarks. Adapting it to the 2-D case is straightforward by disregarding depth ($z$-coordinate), which simply reduces the encoded data volume.
Given $L$ landmarks and $T$ frames, the 2-D version of the Skeleton-DML encoding procedure consists of creating two $L\times T$ matrices (tensors), one for each of the $x$ and $y$ coordinates. These tensors are reshaped so that three consecutive column entries are organized as channels, forming a 3-channel image. The reshaped tensors are then horizontally concatenated, forming a single image. More details are provided in the original work [14] and in [1].
@@ 2.5 Training Protocol
The training protocol follows [1], except for the increasing in the number of epochs (from 20 to 30). At every epoch, we perform on-the-fly data augmentation directly on the landmark sequences prior to their encoding into 2-D images, consisting of rotation, zoom, translation, and horizontal flip to increase data diversity without changing the number of samples. The classifier model is a ResNet-18 initialized with ImageNet weights. Its final convolutional output is first passed through a global average pooling layer, producing a 512-dimensional feature vector. This vector is then fed into two fully connected layers: a 128-unit layer with ReLU activation, followed by an output layer matching the number of classes. The network input is $224\times 224$, while landmark-encoded images have size $126\times(2T/3)$, where $T$ is the number of frames of that particular input video. For the evaluated datasets, $2T/3<224$, so resizing to $224\times 224$ results in upsampling without information loss.
Optimization leverages Adam to minimize cross-entropy, with batch size $64$, learning rate $10^{-4}$, and 30 epochs. Early stopping is triggered if the validation loss fails to decrease for five consecutive epochs. To prevent overfitting, we apply batch normalization after the 512-dimensional layer, dropout (with 50% probability) after the 128-unit layer, and L2 regularization (weight decay of $10^{-4}$) on all trainable parameters. The model selected for evaluation is the checkpoint achieving the highest validation accuracy.
@@ 3 Experimental Methodology
This section outlines the performance assessment, including the Brazilian Sign Language (LIBRAS) datasets, evaluation metrics, experiments, and the computational setup.
@@ 3.1 Datasets
Two popular ISLR LIBRAS datasets were considered in this investigation: MINDS-Libras [18] and LIBRAS-UFOP [5]. In particular, these collections enable direct comparison with [1], whose methodology motivated this study.